X person start walking a 6 miles per hr y person start walking 3 hours later a 10 m/hr how long before they cross paths?

Answer by checkley79(3341)

Assuming they start from the same point & go in the same direction then:
D=RT
D=6T FOR THE FIRST WALKER.
D=10(T-3) FOR THE SECOND WALKER.
SEEING AS THE DISTANCES ARE THE SAME THEN:
6T=10(T-3)
6T=10T-30
6T-10T=-30
-4T=-30
T=-30/-4
T=7.5 HOURS AFTER THE FIRST WALKER LEAVES THEY WILL MEET.
pROOF:
6*7.5=10(7,5-3)
45=10*4.5
45=45